# Towards Scene Graph Anticipation

### v1

@@ Abstract
Spatio-temporal scene graphs represent interactions in a video by decomposing scenes into individual objects and their pair-wise temporal relationships. Long-term anticipation of the fine-grained pair-wise relationships between objects is a challenging problem. To this end, we introduce the task of Scene Graph Anticipation (SGA). We adapt state-of-the-art scene graph generation methods as baselines to anticipate future pair-wise relationships between objects and propose a novel approach SceneSayer. In SceneSayer, we leverage object-centric representations of relationships to reason about the observed video frames and model the evolution of relationships between objects. We take a continuous time perspective and model the latent dynamics of the evolution of object interactions using concepts of NeuralODE and NeuralSDE, respectively. We infer representations of future relationships by solving an Ordinary Differential Equation and a Stochastic Differential Equation, respectively. Extensive experimentation on the Action Genome dataset validates the efficacy of the proposed methods.00footnotetext: $*{}^{*}$hese authors contributed equally to this work.
@@ 1 Introduction
We focus on spatio-temporal scene graphs [15], which is a widely used framework for representing the evolving spatial and temporal relationships among objects. These graphs contain information about the objects present in a video, including their categories, positions, sizes, and spatial dependencies. Simultaneously, they illustrate how these relationships evolve over time, revealing objects’ movement, interactions, and configuration changes across consecutive frames in a video sequence. They facilitate our understanding of dynamic scenes and serve as a valuable tool for addressing downstream tasks in applications such as action recognition and video analysis, where the temporal dynamics of object interactions play a crucial role.
This paper introduces a novel task known as Scene Graph Anticipation (SGA), which, given a video stream, aims to forecast future interactions between objects, as shown in Figure 1. The Scene Graph Anticipation (SGA) task holds significance across diverse domains due to its potential applications and relevance to several downstream tasks. For instance, it contributes to enhanced video understanding by predicting spatiotemporal relationships within video scenes, facilitating improved video analysis and interpreting complex object interactions over time. SGA plays a crucial role in activity recognition, enabling systems to predict future object interactions for more accurate classification and advanced surveillance. Anticipation aids in anomaly detection by identifying deviations from expected object relationships, thereby enhancing the detection of abnormal events in video sequences. Intelligent surveillance systems can also benefit from SGA, allowing systems to predict and respond to security threats by understanding evolving object relationships in monitored environments. Finally, SGA is essential for predicting object movements and interactions for applications in robotics and autonomous systems, contributing to safer and more efficient navigation and decision-making processes.
@@
To tackle the challenges of SGA, we introduce two novel approaches that extend a state-of-the-art scene graph generation method [7]. These approaches utilize object-centric representations of relationships, allowing us to analyze observed video frames and effectively model the dynamic evolution of interactions between objects. This detailed understanding of temporal dynamics is the basis for our proposed methods.
In a departure from traditional sequential modelling, our two approaches adopt a continuous-time perspective. Drawing inspiration from Neural Ordinary Differential Equations (NeuralODE) [3] and Neural Stochastic Differential Equations (NeuralSDE) [17], respectively, we develop methods that capture the latent dynamics governing the evolution of object interactions. By formulating the anticipation problem as solving Ordinary Differential Equations (ODE) and Stochastic Differential Equations (SDE) and thus using a continuous representation of the anticipated relationships, we hope to significantly expand the fidelity of our predictions. We rigorously validated our proposed methods and a strong generation-based baseline on the Action Genome dataset [15], a benchmark for spatio-temporal scene understanding. Our experimental results demonstrate the superior performance of our approaches in accurately anticipating fine-grained pair-wise relationships between objects.
@@ 2.1 Structured Representation
Spatial Graphs. Learning to represent visual content in static data such as 2D/3D images as a spatial graph where the objects act as nodes and edges to describe the visual relation between the objects is called Image Scene Graph Generation (ImgSGG). There has been extensive research in 2D ImgSGG direction following the seminal work of Visual Genome [20]. Following this, the task has been extended to 3D by the introduction of [19]. Recently, a surge of ImgSGG methods explored the role of foundation models in open-vocabulary ImgSGG [5], weakly supervised ImgSGG [18], panoptic ImSGG [42] and zero-shot ImSGG [41].
@@
Spatio-Temporal Graphs. Dynamic visual content, such as videos, provides a more natural set of contextual features describing dynamic interaction between objects. Encoding such content into a structured spatiotemporal graphical representation for frames where nodes describe objects and edges describe the temporal relations is called Video Scene Graph Generation (VidSGG). Early approaches on VidSGG extended previously proposed ImSGG-based methods to the temporal domain. We refer to [43] for a comprehensive survey on earlier work. Recent work explored learning better representations using architectures like Transformers [7, 33] and unbiased representations [27, 16] owing to the long-tailed nature of the benchmark VidSGG datasets Action Genome [15], VidVRD [32].
@@
Applications. Structured representation of visual content has been used in various downstream tasks, including task planning [1], image manipulation [8], visual question answering[6] using scene graphs, video synthesis using action graphs [2] and using scene graph as a knowledge base [21]. To the best of our knowledge, we are the first to formally introduce the task of Scene Graph Anticipation (SGA), propose baseline methodologies, and establish evaluation metrics.  [^2] SGA is related to [38], wherein they use relation forecasting as an intermediate step for action prediction. However, we aim for precise anticipation of future relationships between interacting objects. SGA is also closely related to [26], which constructed a dataset from Action Genome by sampling frames and truncating the videos. However, we differentiate our method by utilizing the complete dataset to train our models. Furthermore, we acknowledge that the approach in [26] resembles the baseline approach proposed in this paper.
2.2 Dynamics Models
@@
Video Prediction. Early video prediction methods treat dynamics modelling as a sequential prediction problem in pixel space using image-level features [22]. Later approaches proposed include using external knowledge as priors [37, 10], better architectural design to model contextual information [40, 30], and focusing on object-centric representations [39]. Recently, there has been a surge in methods proposed that use diffusion models to estimate the distribution of a short future video clip [14, 36]. Contrary to the conventional dense pixel-based generation often seen in video prediction techniques, SGA emphasizes the importance of learning representations that facilitate the prediction of long-term interactions and their nuanced relationships. Our experimental results demonstrate that this approach can effectively estimate these relationship representations for periods extending beyond 30 seconds into the future.
Neural Differential Equations. After the introduction of NeuralODE [3], several methods were proposed to explore latent dynamics models using the framework of NeuralODE. These include methods on trajectory prediction [24], traffic forecasting [25], video generation [28], and multi-agent dynamical systems [12, 29, 13].
@@ 3 Background
3.1 Ordinary Differential Equations (ODEs)
The initial value problem (IVP) for an ODE is given by:
@@ 3.1 Ordinary Differential Equations (ODEs)
Here, $\mathbf{f}:\mathbb{R}^{d}\times\mathbb{R}\rightarrow\mathbb{R}^{d}$ represents a time-varying smooth vector field, and $\mathbf{z}(t)$ is the solution to the IVP. Chen et al. [3] introduced the framework of NeuralODEs wherein they relaxed the time-variance of the vector field $\mathbf{f}$ and parameterized it through neural networks, thus enabling efficient learning of dynamical systems from data. This approach has paved the way for modelling the dynamics of latent states through LatentODEs. Analytical solutions for complex ODEs are typically infeasible; hence, we resort to numerical methods that discretize the time domain into finite intervals, approximating the solution at each step. Thus, they trade off precision with computation time. On the faster, less accurate side, we have single-step methods, such as Euler, while on the slower, more accurate side, we have multistep methods, such as Adams-Bashforth.
3.2 Stochastic Differential Equations (SDEs)
An initial value problem for an SDE is formulated as:
@@ 3.2 Stochastic Differential Equations (SDEs)
Here, $\mu(\mathbf{z}(t),t):\mathbb{R}^{d}\times\mathbb{R}\rightarrow\mathbb{R}^{d}$, $\sigma(\mathbf{z}(t),t):\mathbb{R}^{d}\times\mathbb{R}\rightarrow\mathbb{R}^{(d\times m)}$ represent drift and diffusion terms, respectively and $W_{t}$ denotes $m$-dimensional Wiener process. NeuralSDEs employ neural networks to parameterize drift and diffusion terms, enabling the learning of stochastic processes from data [23, 17]. Solving SDEs analytically is often challenging. Numerical solutions depend on the choice of interpretation of the SDE, which is often connected to the conceptual model underlying the equation. There are many interpretations of an SDE, of which Ito and Stratonovich are popular amongst them. Stratonovich’s interpretation of an SDE is commonly used to model physical systems subjected to noise. SDE solvers use discretization methods to approximate the continuous dynamics of the system over small time intervals. The most common numerical methods for solving SDEs include Euler-Maruyama, Milstein, and Runge-Kutta schemes, each varying in accuracy and computational complexity.
@@ 4 Notation & Problem Description
Given an input video segment $V_{1}^{T}$, we represent it using a set of frames $V_{1}^{T}=\{I^{t}\}_{t=1}^{T}$ defined on discretized time steps $t=\{1,2,\cdots,T\}$, where the total number of observed frames $T$ varies across video segments. A scene graph is a symbolic representation of the objects present in the frame and their pair-wise relationships. In each frame $I^{t}$, we represent the set of objects observed in it using $O^{t}=\{o^{t}_{k}\}_{k=1}^{N(t)}$, where $N(t)$ the total number of objects observed in a frame and varies across frames. Let $\mathcal{C}$ be the set comprising all object categories, then each instance of an object $o_{k}^{t}$ is defined by its bounding box information $\mathbf{b}^{t}_{k}$ and object category $\mathbf{c}^{t}_{k}$, where $\mathbf{b}^{t}_{k}\in[0,1]^{4}$ and $\mathbf{c}^{t}_{k}\in\mathcal{C}$. Let $\mathcal{P}$ be the set comprising all predicate classes that spatio-temporally describe pair-wise relationships between two objects; then each pair of objects $\left(o_{i}^{t},o_{j}^{t}\right)$ may exhibit multiple relationships, defined through predicates $\{p^{t}_{ijk}\}_{k}$ where $p^{t}_{ijk}\in\mathcal{P}$.
We define a relationship instance $r_{ijk}^{t}$ as a triplet $\left(o_{i}^{t},p^{t}_{ijk},o_{j}^{t}\right)$ that combines two distinct objects $\left(o_{i}^{t},o_{j}^{t}\right)$ and a predicate $p^{t}_{ijk}$. Thus, the scene graph $\mathcal{G}^{t}$ is the set of all relationship triplets $\mathcal{G}^{t}=\{r_{ijk}^{t}\}_{ijk}$. Additionally, for each observed object $o_{i}^{t}$ and a pair of objects $\left(o_{i}^{t},o_{j}^{t}\right)$, we use $\hat{\mathbf{c}}_{i}^{t}$ and $\hat{\mathbf{p}}_{ij}^{t}$ to represent the distributions over object categories and predicate classes respectively. Here, $\hat{\mathbf{c}}_{i}^{t}\in\left[0,1\right]^{|\mathcal{C}|}$ and $\hat{\mathbf{p}}_{ij}^{t}\in\left[0,1\right]^{|\mathcal{P}|}$.
@@
4.0.3 Graph Building Strategies.
To build scene graphs for future frames, we employ two strategies that are widely established within VidSGG literature: (a) With Constraint Strategy: This approach enforces a unique interaction constraint between any pair of objects in a scene. Specifically, for a pair of objects $\left(o_{i}^{t},o_{j}^{t}\right)$, there exists a single predicate $p^{t}_{ij}$ that describes the relationship between them, and we incorporate relationship triplets $\{r_{ij}^{t}\}_{ij}$ into the scene graph $\mathcal{G}^{t}$. (b) No Constraint Strategy: In contrast, this method embraces a more complex and detailed graph structure by permitting multiple edges between any pair of objects $\left(o_{i}^{t},o_{j}^{t}\right)$. Specifically, we incorporate all predicted relationship triplets $\{r_{ijk}^{t}\}_{ijk}$ into the scene graph $\mathcal{G}^{t}$.
@@ 5 Technical Approach
Figure 2: Overview of SceneSayer. The forward pass of SceneSayer begins with ORPU, where initial object proposals are generated for each frame. These proposals are then fed to a temporal encoder to ensure that the object representations remain consistent over time. Next, in SCPU, we construct initial relationship representations by concatenating the representations of interacting objects. These representations are further refined using a spatial encoder, embedding the scene’s spatial context into these relationship representations. Then, the representations undergo further enhancement in LDPU, where another temporal encoder fine-tunes them, imbuing the data with comprehensive spatio-temporal scene knowledge. These refined relationship representations from the final observed frame are then input to a Latent Dynamics Model (LDM), where a generative model, either a NeuralODE or a NeuralSDE, generates relationship representations of interacting objects in future frames by solving the corresponding differential equations. Finally, these future relationship representations are decoded to construct anticipated scene graphs.

### v2

@@ Abstract
Spatio-temporal scene graphs represent interactions in a video by decomposing scenes into individual objects and their pair-wise temporal relationships. Long-term anticipation of the fine-grained pair-wise relationships between objects is a challenging problem. To this end, we introduce the task of Scene Graph Anticipation (SGA). We adapt state-of-the-art scene graph generation methods as baselines to anticipate future pair-wise relationships between objects and propose a novel approach SceneSayer. In SceneSayer, we leverage object-centric representations of relationships to reason about the observed video frames and model the evolution of relationships between objects. We take a continuous time perspective and model the latent dynamics of the evolution of object interactions using concepts of NeuralODE and NeuralSDE, respectively. We infer representations of future relationships by solving an Ordinary Differential Equation and a Stochastic Differential Equation, respectively. Extensive experimentation on the Action Genome dataset validates the efficacy of the proposed methods.00footnotetext: denotes equal contribution with names in alphabetical order.
@@ 1 Introduction
We focus on spatio-temporal scene graphs [Ji_2019], which is a widely used framework for representing the evolving spatial and temporal relationships among objects. These graphs contain information about the objects present in a video, including their categories, positions, sizes, and spatial dependencies. Simultaneously, they illustrate how these relationships evolve over time, revealing objects’ movement, interactions, and configuration changes across consecutive frames in a video sequence. They facilitate our understanding of dynamic scenes and serve as a valuable tool for addressing downstream tasks in applications such as action recognition and video analysis, where the temporal dynamics of object interactions play a crucial role.
We introduce a novel task known as Scene Graph Anticipation (SGA), which, given a video stream, aims to forecast future interactions between objects, as shown in Figure 1. The Scene Graph Anticipation (SGA) task holds significance across diverse domains due to its potential applications and relevance to several downstream tasks. For instance, it contributes to enhanced video understanding by predicting spatiotemporal relationships within video scenes, facilitating improved video analysis and interpreting complex object interactions over time. SGA plays a crucial role in activity recognition, enabling systems to predict future object interactions for more accurate classification and advanced surveillance. Anticipation aids in anomaly detection by identifying deviations from expected object relationships, thereby enhancing the detection of abnormal events in video sequences. Intelligent surveillance systems can also benefit from SGA, allowing systems to predict and respond to security threats by understanding evolving object relationships in monitored environments. Finally, SGA is essential for predicting object movements and interactions for applications in robotics and autonomous systems, contributing to safer and more efficient navigation and decision-making processes.
@@
To tackle the challenges of SGA, we introduce two novel approaches that extend a state-of-the-art scene graph generation method [cong_et_al_sttran_2021]. These approaches utilize object-centric representations of relationships, allowing us to analyze observed video frames and effectively model the dynamic evolution of interactions between objects. This detailed understanding of temporal dynamics is the basis for our proposed methods.
In a departure from traditional sequential modelling, our two approaches adopt a continuous-time perspective. Drawing inspiration from Neural Ordinary Differential Equations (NeuralODE) [Chen_2018] and Neural Stochastic Differential Equations (NeuralSDE) [kidger2021efficient], respectively, we develop methods that capture the latent dynamics governing the evolution of object interactions. By formulating the anticipation problem as solving Ordinary Differential Equations (ODE) and Stochastic Differential Equations (SDE) and thus using a continuous representation of the anticipated relationships, we hope to significantly expand the fidelity of our predictions. We rigorously validated our proposed methods and strong generation-based baselines on the Action Genome dataset [Ji_2019], a benchmark for spatio-temporal scene understanding. Our experimental results demonstrate the superior performance of our approaches in accurately anticipating fine-grained pair-wise relationships between objects.
@@ 2 Related Work
Spatial Graphs. Learning to represent visual content in static data such as 2D/3D images as a spatial graph where the objects act as nodes and edges to describe the visual relation between the objects is called Image Scene Graph Generation (ImgSGG). There has been extensive research in 2D ImgSGG direction following the seminal work of Visual Genome [krishna_et_al_visual_2017]. Following this [kim_et_al_3dsgg_2019] extended the task to static 3D (RGB & depth) scene data. Recently, there has been a surge in methods that explored the role of foundation models in ImgSGG task-variants such as open-vocabulary ImgSGG [chen_et_al_expanding_scene_graph_2023], weakly supervised ImgSGG [kim_et_al_llm4sgg_2023], panoptic ImgSGG [zhou_et_al_vlprompt_2023] and zero-shot ImgSGG [zhao_et_al_less_2023, li2023zeroshotvisualrelationdetection].
@@
Spatio-Temporal Graphs. Dynamic visual content, such as videos, provides a more natural set of contextual features describing dynamic interaction between objects. Encoding such content into a structured spatiotemporal graphical representation for frames where nodes describe objects and edges describe the temporal relations is called Video Scene Graph Generation (VidSGG). Early approaches on VidSGG extended previously proposed ImSGG-based methods to the temporal domain. We refer to [zhu_et_al_sgg_2022] for a comprehensive survey on earlier work. Recent work explored learning better representations using architectures like Transformers [cong_et_al_sttran_2021, shit_et_al_relformer_2022] and unbiased representations [nag_et_al_tempura_2023, khandelwal_correlation_2023] owing to the long-tailed datasets Action Genome [Ji_2019], VidVRD [xindi_et_al_vid_vrd_2017].
@@
Applications of Structured Representations. Structured representation of visual content has been used in various downstream tasks, including task planning [agia_et_al_taskography_2022], image manipulation [dhamo_et_al_image_2020], visual question answering[cherian_et_al_vqa_2022], video synthesis using action graphs [bar_et_al_action_2021] and using scene graph as a knowledge base [kurenkov_et_al_sgm_2023]. To the best of our knowledge, we are the first to formally introduce the SGA task and propose baseline methodologies [^2] SGA is related to [Wu_Zhao_Wang_2021], wherein they use relation forecasting as an intermediate step for action prediction. However, we aim for precise anticipation of future relationships between interacting objects. SGA is also closely related to [Mi_2021], which constructed a dataset from Action Genome by sampling frames and truncating the videos. However, we differentiate our method by utilizing the complete dataset to train our models without any pre-processing. Furthermore, we acknowledge that the approach in [Mi_2021] resembles the proposed baseline (Variant-1)..
@@
Video Prediction. Early video prediction methods treat dynamics modelling as a sequential prediction problem in pixel space using image-level features [lee_et_al_stochastic_2018]. Later approaches proposed include using external knowledge as priors [walker2016uncertain, finn2016unsupervised, lu_et_al_mm_prior_kdd], better architectural design to model contextual information [yu2022modular, qi2021learning], and focusing on object-centric representations [wu2023slotformer]. Recently, there has been a surge in methods proposed that use diffusion models to estimate the distribution of a short future video clip [höppe2022diffusion, voleti2022mcvd]. Contrary to the conventional dense pixel-based generation often seen in video prediction techniques, SGA emphasizes the importance of learning relationship representations that facilitate the prediction of the evolution of observed interactions between objects. Our experiments demonstrate that the proposed method can effectively estimate these relationship representations for periods extending beyond 30 seconds into the future.
Neural Differential Equations. Several previous works leveraged the frameworks of NeuralODE [Chen_2018] and NeuralSDE [kidger2021efficient] to explore the learning of latent dynamics models for various tasks of interest such as trajectory prediction [liang_et_al_node_2021], traffic forecasting [liu2023graphbased], video generation [park2021vidode], and simulation of multi-agent dynamical systems [huang2020learning, poli2021graph, zijie_graph_ode_2021].
@@ 3 Background
Ordinary Differential Equations (ODEs) The initial value problem (IVP) is given by:
@@
Here, $\mathbf{f}:\mathbb{R}^{d}\times\mathbb{R}\rightarrow\mathbb{R}^{d}$ represents a time-varying smooth vector field, and $\mathbf{z}(t)$ is the solution to the IVP. Chen et al. [Chen_2018] introduced the framework of NeuralODEs wherein they relaxed the time-variance of the vector field $\mathbf{f}$ and parameterized it through neural networks, thus enabling efficient learning of dynamical systems from data. This approach has paved the way for modelling the dynamics of latent states through LatentODEs. Analytical solutions for complex ODEs are typically infeasible; hence, we resort to numerical methods that discretize the time domain into finite intervals, approximating the solution at each step. Thus, they trade off precision with computation time. On the faster, less accurate side, we have single-step methods, such as Euler, while on the slower, more accurate side, we have multistep methods, such as Adams-Bashforth.
Stochastic Differential Equations (SDEs) An initial value problem is formulated as:
@@
Here, $\mu(\mathbf{z}(t),t):\mathbb{R}^{d}\times\mathbb{R}\rightarrow\mathbb{R}^{d}$, $\sigma(\mathbf{z}(t),t):\mathbb{R}^{d}\times\mathbb{R}\rightarrow\mathbb{R}^{(d\times m)}$ represent drift and diffusion terms, respectively and $\mathbf{W}(t)$ denotes an $m$-dimensional Wiener process. NeuralSDEs employ neural networks to parameterize both drift and diffusion terms thus enabling the learning of stochastic processes from data [Li_2021, kidger2021efficient]. Solving SDEs analytically is often challenging and infeasible. Numerical solutions depend on the choice of interpretation of the SDE, which is often connected to the conceptual model underlying the equation. There are many interpretations of an SDE, of which Stratonovich is one of the most popular [oksendal03]. Stratonovich’s interpretation of an SDE is commonly used to model physical systems subjected to noise. SDE solvers use discretization methods to approximate the continuous dynamics of the system over small time intervals. The most common numerical methods for solving SDEs include Euler-Maruyama, Milstein, and Runge-Kutta schemes, each varying in accuracy and computational complexity.
@@ 4 Notation & Problem Description
Given an input video segment $V_{1}^{T}$, we represent it using a set of frames $V_{1}^{T}=\{I^{t}\}_{t=1}^{T}$ defined on discretized time steps $t=\{1,2,\cdots,T\}$, where the total number of observed frames $T$ varies across video segments. A scene graph is a symbolic representation of the objects present in the frame and their pair-wise relationships. In each frame $I^{t}$, we represent the set of objects observed in it using $O^{t}=\{o^{t}_{k}\}_{k=1}^{N(t)}$, where $N(t)$ denotes the total number of objects observed in $I^{t}$. Let $\mathcal{C}$ be the set comprising all object categories, then each instance of an object $o_{k}^{t}$ is defined by its bounding box information $\mathbf{b}^{t}_{k}$ and object category $\mathbf{c}^{t}_{k}$, where $\mathbf{b}^{t}_{k}\in[0,1]^{4}$ and $\mathbf{c}^{t}_{k}\in\mathcal{C}$. Let $\mathcal{P}$ be the set comprising all predicate classes that spatio-temporally describe pair-wise relationships between two objects; then each pair of objects $\left(o_{i}^{t},o_{j}^{t}\right)$ may exhibit multiple relationships, defined through predicates $\{p^{t}_{ijk}\}_{k}$ where $p^{t}_{ijk}\in\mathcal{P}$. We define a relationship instance $r_{ijk}^{t}$ as a triplet $\left(o_{i}^{t},p^{t}_{ijk},o_{j}^{t}\right)$ that combines two distinct objects $\left(o_{i}^{t},o_{j}^{t}\right)$ and a predicate $p^{t}_{ijk}$. Thus, the scene graph $\mathcal{G}^{t}$ is the set of all relationship triplets $\mathcal{G}^{t}=\{r_{ijk}^{t}\}_{ijk}$. Additionally, for each observed object $o_{i}^{t}$ and a pair of objects $\left(o_{i}^{t},o_{j}^{t}\right)$, we use $\hat{\mathbf{c}}_{i}^{t}\in\left[0,1\right]^{|\mathcal{C}|}$ and $\hat{\mathbf{p}}_{ij}^{t}\in\left[0,1\right]^{|\mathcal{P}|}$ to represent the distributions over object categories and predicate classes respectively. Here $\sum_{k}\hat{{c}}_{ik}^{t}=1$, $\sum_{k}\hat{{p}}_{ijk}^{t}=1$.
@@
Graph Building Strategies. To build scene graphs for future frames, we employ two strategies that are widely established within the VidSGG literature:
  • With Constraint Strategy: This approach enforces a unique interaction constraint between any pair of objects in a scene. Specifically, for a pair of objects $\left(o_{i}^{t},o_{j}^{t}\right)$, there exists a single relationship predicate $p^{t}_{ij}$ that describes the interaction between the objects, and we incorporate relationship triplets $\{r_{ij}^{t}\}_{ij}$ into the scene graph $\mathcal{G}^{t}$.
  • No Constraint Strategy: In contrast, this approach embraces a more complex and detailed graph structure by permitting multiple edges (where each edge describes a relationship predicate) between any pair of interacting objects $\left(o_{i}^{t},o_{j}^{t}\right)$. Specifically, we incorporate all predicted relationship triplets $\{r_{ijk}^{t}\}_{ijk}$ into the scene graph $\mathcal{G}^{t}$.
@@ 5 Technical Approach
Figure 2: Overview of SceneSayer. The forward pass of SceneSayer begins with ORPU, where initial object proposals are generated for each frame. These proposals are then fed to a temporal encoder to ensure that the object representations remain consistent over time. Next, in SCPU, we construct initial relationship representations by concatenating the representations of interacting objects. These representations are further refined using a spatial encoder, embedding the scene’s spatial context into these relationship representations. Then, the representations undergo further enhancement in LDPU, where another temporal encoder fine-tunes them, imbuing the data with comprehensive spatio-temporal scene knowledge. These refined relationship representations from the final observed frame are then input to a Latent Dynamics Model (LDM), where a generative model, either a NeuralODE or a NeuralSDE, generates relationship representations of interacting objects in future frames by solving the corresponding differential equations. Finally, these future representations are decoded into relationship predicates to construct anticipated scene graphs.